OpenStudy (anonymous):

Which of the following represents the linear equation 2x = y + 3 in slope−intercept form?

OpenStudy (anonymous):

x + y = 3 2x − y = 3 y = 2x − 3 y = 2x + 3

OpenStudy (radar):

Only the last two are in the proper form, you have a 50-50 chance now.

OpenStudy (anonymous):

its c correct? I'm just double checking my answer

OpenStudy (radar):

Yes.

OpenStudy (anonymous):

thanks